use the distance formula to find the length of the line segment between the pair of points. Round to hundredths.
(5, 4) and (-1, -3)

Answer by dabanfield(803)

d^2 = (x1-x2)^2 + (y1-y2)^2 where d is the distance between the two points (x1,y1) an d(x2,y2).
So in this case
d^2 = (5-(-1))^2 + (4-(-3))^2
d^2 = 6^2 + 7^2
d^2 = 36 + 49 = 85
d = ?